# Multi-modal Gaussian Process Variational Autoencoders for Neural and Behavioral Data

### v1

@@ 4.2 Application to fly experimental data
Next we look to evaluate the MM-GPVAE on a real-world multi-modal dataset. Here, we consider a whole-brain calcium imaging from an adult, behaving Drosophila [35]. We isolate 1000 calcium traces recorded using SCAPE microscopy [36] from an animal while it performs a variety of distinct behaviors fixed on a spherical treadmill. These 1000 calcium traces are dispersed widely and uniformly across the central fly brain and were randomly picked from the dataset to use in our model. Alongside the neural measurements, eight 2-d limb positions are extracted from a recorded video using the software tool DeepLabCut [37]. These simultaneously recorded calcium traces and limb position measurements are split into 318 trials of 35 time-bins sampled at 70hz. Each trial has one of 5 corresponding behavioral labels (still, running, front grooming, back grooming and abdomen bending) determined via a semi-supervised approach from the tracked limb position measurements [38]. Importantly, these behavioral labels are not used when fitting the MM-GPVAE. For additional information on how we isolate this behavioral data and neural traces, see the supplemental materials.
For this set-up, we consider a variant of the MM-GPVAE where we have removed the non-linearity for the neural data modality, and instead consider Gaussian observations with an additional parameter controlling the observation variance for the fluorescence traces (akin to the original formulation of GPFA [30]). We do this because calcium imaging traces take on continuous values and are not well-described by Poisson observations, though in principle other observation likelihoods could be used here [39]. Next, we consider a 7-dimensional latent behavioral subspace and 26 dimensional neural subspace, with 5 of these dimensions being shared across modalities. These choices were made through initial exploration of the model and examining cross-validated model performance, though the results we show are robust to a wide range of dimensionality choices for each of the subspaces.
@@
The MM-GPVAE is able to successfully reconstruct both behavioral trajectories and calcium traces in held-out trials for these data. The top of Figure 4(c) shows the true and decoded 16 limb position measurements with 5 highlighted for clarity. The bottom part of figure 4(c) shows 6 randomly selected calcium traces alongside their model reconstructions. In each case, the MM-GPVAE can roughly capture the temporal trends in this multi-modal dataset. To analyze the shared and independent latent subspaces of these data we consider a 2d projection of each latent space, and plot the mean latent value in that subspace calculated each trial. We additionally color-code the trial according to the behavioral label. In these shared and independent latent subspaces, we find that the "still" behavioral state is well separated from many of the other behavioral conditions in the neural and shared subspaces. Here, we show two others (bending and running) as an illustration (Figure 4(d)). We refer the reader to the supplement for visualizing all 5 behaviors in the latent space. Lastly, we illustrate how much each subspace contributes to the overall variability of neural and behavioral data. Figure 4(b) shows the variance of the neural (left) and behavioral (right) data calculated for each trial where the reconstruction of the data is generated either from the shared subspace, or either of the independent subspaces. Here, we find that the shared subspace for each modalities contributes more to the overall variability in the data than either of the independent subspaces, suggesting a large fraction of shared variability between the data modalities.
@@ 4.3 Application to moth experimental data
Next, we evaluate the MM-GPVAE on a dataset of a hawkmoth (Manduca sexta) tracking a moving flower. The hawkmoth is an agile flier that is able to closely follow swiftly moving targets while hovering in midair, making it a model organism for the study of sensorimotor control [40]. The modalities in our dataset consist of a time-series of images of a visual stimulus temporally paired with electromyography (EMG) signals recorded during 20-second experimental trials. Each image is an event-based reconstruction (inspired by insect visual system) of a white floral target moving laterally during a trial as viewed in the hawkmoth’s frame of reference [41] (Fig 5(a)). The floral target moves sinusoidally at 1 Hz against a black background. Each pixel in a video frame can occupy a polarity of -1, 0 or +1 based on whether its luminosity has respectively decreased, unchanged or increased as compared to the previous frame in the temporal sequence. Temporally paired with the floral stimulus, EMG recordings are precisely measured spike trains from the 10 major flight muscles of the hawkmoth responding to the floral target while flying on a tether and flapping at about 22 Hz. Altogether, the 10-muscle recordings form a near-complete motor program that controls its flight during target tracking response [42].
@@
We fit the MM-GPVAE model to these data using a 2-dimensional latent space to describe the moving visual stimulus and a 3-dimensional latent space to describe the Poisson rates from the 10 motor units. Of these, one latent dimension is shared between modalities. Here, a small number of dimensions are used both because only a few dimensions were needed to be able to accurately reconstruct the data of each modality, and also because it allows for easier interpretation of the latent dynamics with respect to the experiment. We split a 20 second recording into 6 trials of equal length for model fitting. Here, it is worth noting that the trials are much longer ( 3300 timebins) than in the previous experiments. To encourage the visual-stimulus latent space to capture the slow dynamics of the moving flower, and the neural latent space to capture the fast dynamics of the flapping wings, we initialized the length-scale () parameter of the GP kernel at a value of 300 for the image latent, 10 for the neural latent, and 150 for the latent component shared between modalities. We find that the MM-GPVAE is able to accurately reconstruct both the neural rates (Fig 5(b)) and visual stimulus (Fig 5(c)) from its low dimensional latent space. We see that the neural rates are strongly modulated by the wing-flap oscillation at approximately 22 Hz. This is of course also seen in the neural latent space identified by the MM-GPVAE. At the bottom of Figure 5(b) we see that the first PC of the 2-dimensional latent space closely aligns with the torque measurement along the z-axis of the moth motion, a measurement previously identified to be closely associated with wing beating [42]. The MM-GPVAE is also able to reconstruct the visual stimulus data, as seen in 5(c). When viewing the latent space corresponding to the visual stimulus modality, we find that the value of the smoothly evolving latent variable closely tracks the position of the stimulus in x dimension (the only dimension along which it moves, see Figure 5(e)). Finally, the shared latent dimension shows some modulation with wing-flapping, but exhibits temporal structure not obviously related to the flapping frequency or stimulus motion. This may correspond to a longer-timescale motor behavior such as the variation in how strongly the hawkmoth is responding to track the stimulus, which could depend on the degree of its attention to the stimulus and its state of motivation. Importantly, we find that including a shared dimension better reconstructs the data than not including one, suggesting that there is a component of neural dynamics shared with stimulus dynamics. Specifically, we find that if we remove the shared latent we see approximately a 70% drop in mse in reconstruction of the image modality and about a 5% drop for reconstruction of the rates. Because the neural rates have such a strong wing-flapping modulation, we expect the slower-moving dynamics to account for only a small amount of the variability here.
Lastly, to emphasize the added interpretability we get from the linear layer in the neural likelihood (GPFA), we plot the weights and that map the shared and independent neural latents to the log-rate of the hawkmoth neurons. This is easiest to visualize in this example dataset because there is only a small neural population from which we record spikes. Figure 5 (d) shows each latent loading on to each of the ten neurons. Here, we z-score the weights along each dimension for easy visualization. We see that certain neurons are more heavily driven by the neural-only latents whereas others are more strongly modulated by the shared latent component. Future analysis of these data may help us further uncover how these shared and independent latent dynamics relate to the hawkmoth tracking and specific sensorimotor neurons and function in this experiment.
5 Comparison to Existing Approaches
Our work contrasts in important ways with similar recently-developed multi-modal latent variable models for neural time series experiments. One closely related model, Targeted Neural Dynamical Modeling (TNDM) [11], was specifically designed to non-linearly separate the neural latent representation into behaviorally relevant and irrelevant subspaces. However, this work was evaluated on a dataset with a relatively simple behavioral paradigm, and fails to to reconstruct complex behavioral modalities (so is not well suited to reconstructing stimuli, for example). Another related multi-modal time-series neuroscience model, Preferential Subspace Identification (PSID) [12], was designed to linearly separate the neural latent representation into behaviorally relevant and irrelevant subspaces. However, PSID is restricted in its ability to describe more complex dynamics, and similarly to TNDM only performs well when behavioral or stimulus modalities have simple structure due to lack of a DNN decoder. To demonstrate the capabilities of MM-GPVAE alongside these completing methods, we evaluate the MM-GPVAE, TNDM, and PSID on the simple dataset used in TNDM paper as well as our simulated rotating and scaling MNIST example.
We first compare our MM-GPVAE with TNDM [11] and PSID [12] using our simulated data of rotating and scaling MNIST digit ‘3’ and simultaneous Poisson neural rates to show the reconstruction performance of our model in comparison to these two competing models. Error on both data modalities as well as reconstructions of each modality on held-out trials can be found in the supplement.
To further compare these models in a real-world setting, we then implement PSID, TNDM, and MMGPVAE in a simple neuroscience example where all models are capable of recovering behavioral trajectories as well as neural rates. We use the primate reaching data used in the TNDM paper. Here, the 2D reaching task is simple, and so linear dynamics as well as our deep neural network can reconstruct both modalities well. The results from this can be seen in Figure 6. While both TNDM and PSID can provide only two subspaces in which only one can separate the 8 reaching positions (relevant subspace), MM-GPVAE can provide visualization for 3 subspaces, 2 modality specific, and 1 shared, and we can see a clear separation with both the behavioral-only subspace as well as shared subspace.
@@ 6 Conclusion
In this work, we have introduced the multi-modal Gaussian Process variational autoencoder (MM-GPVAE) to identify temporally evolving latent variables for jointly recorded neural activity and behavioral or stimulus measurements. We parameterized the model in the Fourier domain to better extract identifiable temporal structure from high-dimensional time-series data. We first demonstrated that the Fourier-mean field representation is able to better recover the true latent variables in a single-modality setting compared to previous approaches in the time domain. We then show that our multi-modal GPVAE can also accurately recover latent structure across data modalities, which we demonstrate using a smoothly rotating and scaling MNIST digit alongside simulated neural spike trains. We then show that the model can be flexibly adapted to multiple real-world multi-modal behavioral settings. We implement the MM-GPVAE on calcium imaging traces and tracked limb positions of Drosophila and visualize the latent embeddings of distinct fly behaviors in shared and independent latent subspaces. We also use the MM-GPVAE on spikes from hawkmoth flight muscles as the animal tracks an oscillating visual stimulus. Here, our model extracted latent variables from each modality that corresponds to distinct experimentally relevant variables - wing flapping and a much slower stimulus movement. Here, we also demonstrated that the MM-GPVAE can be adapted to long time-series data to identify latent variables at widely varying timescales.
Choice of prior: In this work we focus on the use of GP priors to describe dynamics, though many other choices could have been made. As stated above, the existing multi-modal time-series models in neuroscience using linear dynamics and RNNs have limitations when compared with the MM-GPVAE. However, other dynamical approaches such as switching linear dynamical systems or non-linear dynamics models exist in neuroscience and are powerful in single modality settings [9, 43]. Such models contain dependencies across the latent dynamical variables, and so require reformulation and developing new inference to adapt them to multi-modal settings, and do not afford the interpretability and scalability gains seen with the Fourier approach. Thus, development of multi-modal versions of different dynamical models and evaluating their strengths and limitations alongside the MM-GPVAE remains an important avenue of future work in neuroscience.
Limitations: One limitation of our model is that it is specifically designed for temporally aligned time-series data, therefore requiring the data from both modalities to be recorded at matching time intervals or needing resampling to align them (see supplement). However, techniques such as time warping can be implemented to address this limitation. The use of the GP prior with RBF kernel and pruned Fourier representation also assumes latent dynamics are varying smoothly in time, precluding the appropriateness of our model for data with abrupt state-changes. Lastly, despite the improvements of the Fourier domain representation in our model, model performance can be sensitive to optimization step-size, covariance parameter , as well as initial length-scale parameter values, pruning thresholds and marginal variance . See supplement for more information and example code.
Broader impact: Our work provides a general model for exploratory data analysis for multi-modal time-series data. Though we develop the model with neuroscience experiments in mind, in principle the model could be adapted to other settings where the assumption of smoothly evolving latent dynamics holds across distinct data modalities. We do not anticipate any potential negative societal impacts of our work.
@@
Additional notes on inference: All models in this manuscript were trained end-to-end in PyTorch using the ADAM optimizer. Training was done on a Macbook Pro with Apple M1 max chip and all evaluations took less than an hour to fit. All encoder and decoder neural networks were standard feedforward neural networks with ELU activation functions. Additional details on neural network architectures for the experiments can be found in the supplement. The code for MM-GPVAE will be available upon formal publication.
Supplementary Material
Additional comparison with Targeted Neural Dynamical Modeling (TNDM) and Preferential Subspace Identification (PSID)
@@
In order to compare MM-GPVAE with Targeted Neural Dynamical Modeling (TNDM) [11] and Preferential Subspace Identification (PSID) [12], we evaluated both TNDM and PSID with our rotating and scaling MNIST digit ’3’ alongside Poisson neural rates. In Figure 7, we show the reconstruction of both the behavioral modality, the MNIST digit ’3’, and Poisson neural rates for each model. In addition to these, we also provide the error on both data modalities on held-out trials. Both TNDM and PSID have linear decoders to observations, thus are not capable of accurately describing the rotating digit. On top of this, we see that the MMGPVAE shows better reconstruction of the underlying smooth neural rates.
@@
To provide a more complete picture of the ability of the MM-GPVAE to both reconstruct simulated ’behavioral’ and neural data, and to accurately recover the true underlying latent variables, we show additional performance evaluations here. Figure 8 shows an example of 24 reconstructed 3s from the evaluation shown in Figure 3 of the manuscript. Again, here we reconstruct a scaling and rotating MNIST digit ’3’ alongside 100 neural spike trains, where one latent dimension is shared across modalities. Figure 8 also shows latent reconstructions and 5 example neural rates in 5 held out trials. Using the same set-up, we additionally run the MM-GPVAE with MNIST digit ’2’, and show an example of 24 reconstructed images as well as 3 example neural rates (of 100) on 5 held-out trials in Figure 9.
In addition to these, we also extend figure 3(a) from our manuscript with all the latent trajectories for our simulated data comparisons, including estimates from GPFA and GPVAE. We additionally provide clarification for the ‘shared’ latent for our single modality comparisons (GP-VAE and GPFA). Because in GPFA and GPVAE there is ambiguity as to which latent is “shared” (as these are unimodal models), attached Figure 10 panels (b) and (c) show the mean squared error for both of the possible latents for GPVAE and GPFA. Here, we note that the MMGPVAE is able to ‘disentangle’ the shared representation across the modalities, something that each unimodal model is unable to do. We visualize both the average mean-squared error (c) on held-out data as well as a per-trial mean-squared error scatter-plot (b) to provide a better sense of model performance.
@@
For the MM-GPVAE, both the simulated and real-world multi-modal evaluations used similar neural net architectures. However, there were some modifications of the nodes/layers that were unique to each evaluation. This was necessary as the latent dimensionality was different across our different evaluations, and certain behavioral reconstructions, especially those that were high-dimensional, required richer neural network parameterizations. The schematic for the MM-GPVAE neural network architecture in our simulated example can be found in Figure 11, and the schematics for the neural network architectures for the real-world multi-modal datasets can be found in Figure 12 (fly) and Figure 13 (hawkmoth). For our evaluation on the dataset used in [11], whose results are above, the architecture can be found in Figure 14. Note that for all experiments, each modality is encoded to its own set of variational means and variances (transformed into the Fourier domain). The encoded means and variances representing the shared latents are then summed to give the encoded shared latents means and variances. Across all evaluations, we parameterized our latents in the Fourier domain and converted back to the time domain before decoding.
@@
Pre-processing: We isolated 1000 raw calcium traces from [35] for use with our model for the evaluations in section 4.2 of the manuscript. This dataset also contained x,y positions from 8 tracked limbs positions [37]. The data came as one continuous recording, and at every time point there was a behavioral categorization probability for one of 6 distinct behaviors (undefined, still, running, front grooming, back grooming, abdomen bending) determined via the algorithm outlined in [38]. To split this continuous recording into trials, we found segments of the recording where, for 35 continuous samples, a single behavior was estimated at a probability. This generated 318 total trials where each trial was of one of 5 possible behaviors. There was no section of the recording where there was an ’undefined’ behavior for 35 continuous samples, so there are no undefined behavioral trials in our analysis. To avoid the possibility of shared latent to contain dynamics that primarily drive one modality, we use a specific cross-validation procedure. Specifically, after first identifying the appropriate number of dimensions per modality, we increased the number of shared dimensions sequentially until predictive performance no longer improves, assuring we had no more shared dimensions than needed. Additionally, this is more likely to be the case when the dimensionality of the shared subspace is too large.
Additional evaluation: Though we show the time-course of the behavioral reconstruction in the main manuscript, we show a 2-d reconstruction on 2 example trials in Figure 15. Here we plot true x,y positions for 2 trails for each of the 8 limb positions alongside our model’s reconstruction. We can see here that our model also captures the spatial information in the behavioral data. We also show for comparison the 2-d depiction of the shared and independent latent representation of all trials in the dataset (Figure 16), with all five of the behaviors labelled. You can see here that the ’still’ behavior separates well from all other behaviors in the shared and neural subspaces.

### v2

@@ 4.2 Application to fly experimental data
Next, we look to evaluate the MM-GPVAE on a real-world multi-modal dataset. Here, we consider a whole-brain calcium imaging from an adult, behaving Drosophila (Schaffer et al., 2021). We isolate 1000 calcium traces recorded using SCAPE microscopy (Voleti et al., 2019) from an animal while it performs a variety of distinct behaviors fixed on a spherical treadmill. These 1000 calcium traces are dispersed widely and uniformly across the central fly brain and were randomly picked from the dataset to use in our model. Alongside the neural measurements, eight 2-D limb positions are extracted from a recorded video using the software tool DeepLabCut (Mathis et al., 2018). These simultaneously recorded calcium traces and limb position measurements are split into 318 trials of 35 time-bins sampled at 70hz. Each trial has one of 5 corresponding behavioral labels (still, running, front grooming, back grooming and abdomen bending) determined via a semi-supervised approach from the tracked limb position measurements (Whiteway et al., 2021). Importantly, these behavioral labels are not used when fitting the MM-GPVAE. For additional information on how we isolate this behavioral data and neural traces, see the appendix. We consider a 7-dimensional latent behavioral subspace and 26-dimensional neural subspace, with 5 of these dimensions being shared across modalities. These choices were made through initial exploration of the model and examining cross-validated model performance (see appendix), though the results we show are robust to a wide range of dimensionality choices for each of the subspaces.
@@
The MM-GPVAE is able to successfully reconstruct both behavioral trajectories and calcium traces in held-out trials for these data. The top of Figure 4(c) shows the true and decoded 16 limb position measurements with 3 highlighted for clarity. The bottom part of Figure 4(c) shows 6 randomly selected calcium traces alongside their model reconstructions. In each case, the MM-GPVAE can roughly capture the temporal trends in this multi-modal dataset. To analyze the shared and independent latent subspaces of these data we consider a 2d projection of each latent space, and plot the mean latent value in that subspace calculated each trial. We additionally color-code the trial according to the behavioral label. In these shared and independent latent subspaces, we find that the "still" behavioral state is well separated from many of the other behavioral conditions in the neural and shared subspaces. Here, we show two others (bending and running) as an illustration (Fig 4(d)). (For visualizing all 5 behaviors in the latent space, see Figure 19). Lastly, we illustrate how much each subspace contributes to the overall variability of neural and behavioral data. Figure 4(b) shows the variance of the neural (left) and behavioral (right) data calculated for each trial where the reconstruction of the data is generated either from the shared subspace, or either of the independent subspaces. We find that the shared subspace for each modality contributes more to the overall variability in the data than either of the independent subspaces, suggesting a large fraction of shared variability between the data modalities.
@@ 4.3 Application to moth experimental data
Next, we evaluate the MM-GPVAE on a dataset of a hawkmoth (Manduca sexta) tracking a moving flower. The hawkmoth is an agile flier that is able to closely follow swiftly moving targets while hovering in midair, making it a model organism for the study of sensorimotor control (Sprayberry and Daniel, 2007). The modalities in our dataset consist of a time series of images of a visual stimulus temporally paired with electromyography (EMG) signals recorded during 20-second experimental trials. Each image is an event-based reconstruction (inspired by insect visual system) of a white floral target moving laterally during a trial as viewed in the hawkmoth’s frame of reference (Sikandar et al., 2023) (Fig 5(a)). The floral target moves sinusoidally at 1 Hz against a black background. Each pixel in a video frame can occupy a polarity of -1, 0 or +1 based on whether its luminosity has respectively decreased, unchanged or increased as compared to the previous frame in the temporal sequence. Temporally paired with the floral stimulus, EMG recordings are precisely measured spike trains from the 10 major flight muscles of the hawkmoth responding to the floral target while flying on a tether and flapping at about 22 Hz. Altogether, the 10-muscle recordings form a near-complete motor program that controls its flight during target tracking response (Putney et al., 2019).
We fit the MM-GPVAE model to these data using a 2-dimensional latent space to describe the moving visual stimulus and a 2-dimensional latent space to describe the Poisson rates from the 10 motor units. Of these, one latent dimension is shared between modalities. Here, a small number of dimensions are used because only a few dimensions were needed to be able to accurately reconstruct the data of each modality (for more information about dimensionality selection, see appendix).
@@
We find that the MM-GPVAE is able to accurately reconstruct both the neural rates (Fig 5(b)) and visual stimulus (Fig 5(c)) from its low dimensional latent space. We see that the neural rates are strongly modulated by the wing-flap oscillation at approximately 22 Hz. This is also seen in the neural latent space identified by the MM-GPVAE. At the bottom of Figure 5(b) we see that the 1-dimensional latent space closely aligns with the torque measurement along the z-axis of the moth motion, a measurement previously identified to be closely associated with wing beating (Putney et al., 2019). The MM-GPVAE is also able to reconstruct the visual stimulus data, as seen in 5(c). When viewing the latent space corresponding to the visual stimulus modality, we find that the value of the smoothly evolving latent variable closely tracks the position of the stimulus in x dimension (the only dimension along which it moves, see Fig 5(e)). Finally, the shared latent dimension shows some modulation with wing-flapping but exhibits temporal structure not obviously related to the flapping frequency or stimulus motion. This may correspond to a longer-timescale motor behavior such as the variation in how strongly the hawkmoth is responding to track the stimulus, which could depend on the degree of its attention to the stimulus and its state of motivation. Importantly, we find that including a shared dimension better reconstructs the data than not including one, suggesting that there is a component of neural dynamics shared with stimulus dynamics. Specifically, we find that if we remove the shared latent we see approximately a 70% reduction the reconstruction accuracy of the image modality and about a 5% reduction in the reconstruction of the rates (Fig 21). Because the neural rates have a strong wing-flapping modulation, we expect the slower-moving dynamics to account for only a small amount of the variability here. Lastly, to demonstrate the interpretability advantage of linear decoder layer used for the neural likelihood, we visualize how each of the 10 motor neurons is loaded onto either the neural-only dynamics or the dynamics shared with the moving stimulus Figure 5(d). For a closer look a individual neurons modulations by shared and independent latent components, see Figure 21(b).
@@ 5 Conclusion
In this work, we have introduced the multi-modal Gaussian Process variational autoencoder (MM-GPVAE) to identify temporally evolving latent variables for jointly recorded neural activity and behavioral or stimulus measurements. We first parameterize the model in the Fourier domain to better extract identifiable temporal structure from high-dimensional time-series data. We then combine a linear decoder for neural data with a deep neural network decoder for behavioral or stimuli data to enhance the applicability of the MM-GPVAE across experimental settings while retaining the ability to identify neural tuning with respect to shared and independent temporally evolving latents.We show that our multi-modal GPVAE can accurately recover latent structure across data modalities, which we demonstrate using a smoothly rotating and scaling MNIST digit alongside simulated neural spike trains. We then show that the MM-GPVAE can be flexibly adapted to multiple real-world multi-modal settings. First, we analyze calcium imaging traces and tracked limb positions of fruit flies (Drosophila) exhibiting various distinct behaviors. Then, we apply the model to spike trains from hawkmoth flight muscles as well as the oscillating visual stimulus the animal tracks during flight.
Comparison to multi-modal approaches in neuroscience: Our work contrasts in important ways with two similar recently-developed multi-modal latent variable models for time series data in neuroscience. One closely related model, Targeted Neural Dynamical Modeling (TNDM) (Hurwitz et al., 2021), was specifically designed to non-linearly separate the neural latent representation into behaviorally relevant and irrelevant subspaces using a latent space governed by a recurrent neural network (RNN). However, this work was evaluated on a dataset with a relatively simple behavioral paradigm, and fails to accurately model the behavioral modality when it is higher dimensional or more complex (Fig 6). Another related multi-modal time-series neuroscience model, Preferential Subspace Identification (PSID) (Sani et al., 2021), linearly separates the neural latent representation into behaviorally relevant and irrelevant subspaces. However, PSID may be limited in that uses latent dynamics are governed by a linear state-space model, and, similar to TNDM, is restricted in capturing complex behavioral modalities well due a linear decoder to behavioral data (Fig 6). For a more thorough evaluation and discussion to these approaches, see the appendix.
Choice of prior: In this work we focus on the use of GP priors to describe latent dynamics, though other temporal dependencies are of course possible. Though TNDM and PSID have certain limitations when compared with the MM-GPVAE (Fig 6), these are not necessarily due to their latent temporal prescriptions. These models could be modified with more flexible decoding schemes to better isolate the role of the different latent dynamics in characterizing shared and independent subspaces compared to the MM-GPVAE. In addition, other choices such as switching linear or non-linear latent dynamics models are powerful in single modality settings (Hernandez et al., 2018; Linderman et al., 2017; Glaser et al., 2020) and could also be adapted to multi-modal settings. Our choice of a GP prior specifically is well suited to experimental paradigms where the latent dynamics are hypothesized to be smooth. Other approaches might be more appropriate when this smoothness assumption is unlikely to hold. Thus, development of multi-modal versions of different latent dynamical models in neuroscience and evaluating their strengths and limitations alongside the MM-GPVAE remains an important avenue of future work in neuroscience.
6 Reproducibility Statement
@@
All models in this manuscript were trained end-to-end in PyTorch using the ADAM optimizer. Training was done on a Macbook Pro with Apple M1 max chip and all evaluations took less than an hour to fit. All encoder and decoder neural networks were standard feedforward neural networks with ELU activation functions. Additional details on neural network architectures for the experiments can be found in the appendix. An implementation of MM-GPVAE can be found at: GitHub Repository for MM-GPVAE.
7 Ethics Statement
Our work provides a general model for exploratory data analysis for multi-modal time-series data. Though we develop the model with neuroscience experiments in mind, in principle the model could be adapted to other settings where the assumption of smoothly evolving latent dynamics holds across distinct data modalities. We do not anticipate any potential negative societal impacts of our work.
Appendix
Contrasting the MM-GPVAE with existing approaches
@@ 7 Ethics Statement
We evaluate our MM-GPVAE model alongside both Targeted Neural Dynamical Modeling (TNDM) (Hurwitz et al., 2021) and Preferential Subspace Identification (PSID) (Sani et al., 2021) using our synthetic data of a rotating and scaling MNIST digit ‘3’ and corresponding Poisson neural rates. In Figure 6, we show the reconstruction of both the behavioral modality, the MNIST digit ‘3’, and Poisson neural rates for each model. In addition to these, we also provide the error of the reconstruction performance on both data modalities on held-out trials. We find that the MM-GPVAE is able to better reconstruct data from both modalities in this simulated example than the competing approaches.
Next, to compare these models in a real-world setting, we implement PSID, TNDM, and MM-GPVAE on data from a simpler neuroscience experiment where all models are capable of recovering both behavioral trajectories as well as neural rates. We focus on a simpler behavioral task because these competing approaches are not well-suited to reconstruct complex behavioral modalities due to a lack of neural network decoder. We use the primate reaching data from Hurwitz et al. (2021) for these evaluations. Here, a linear decoder as well as our deep neural network can reconstruct both modalities well. The visualizations of the latent spaces for these data can be seen in Figure 7. While both TNDM and PSID can provide only two subspaces in which only one can separate the 8 reaching positions (relevant subspace), MM-GPVAE can provide visualization for 3 subspaces, 2 modality specific, and 1 shared, and we can see a clear separation with both the behavioral-only subspace as well as shared subspace.
@@
To provide a more complete picture of the ability of the MM-GPVAE to both reconstruct simulated behavioral and neural data, and to accurately recover the true underlying latent trajectories, we show additional performance evaluations here in a simulated setting. Figure 8 shows an example of 24 reconstructed 3s from the evaluation shown in Figure 3 of the manuscript. Again, here we reconstruct a scaling and rotating MNIST digit ‘3’ alongside 100 neural spike trains, where one latent dimension is shared across modalities. Figure 8 also shows latent reconstructions and 5 example neural rates in 5 held out trials. Using the same set-up, we additionally run the MM-GPVAE with MNIST digit ‘2’, and show an example of 24 reconstructed images as well as 3 example neural rates (of 100 neurons) on 5 held-out trials in Figure 9.
In addition to these, we extend Figure 3(a) from our manuscript with all the latent trajectories for our simulated data comparisons. Because in our unimodal (GPFA and GP-VAE) comparisons, there is ambiguity as to which latent is “shared”, we simply select the latent trajectory which most closely matches the true underlying latent variable for each comparison. Panels (b) and (c) in Figure 10 show the reconstruction mean squared error of the MM-GPVAE compared to the unimodal models GP-VAE and GPFA. Here, we include a version of each of the unimodal models with two different encoder representations, one which represents what the GPFA and MM-GPVAE would typically have access to - simply the data of their own modality (i.e. neural data for GPFA, image/behavioral data for GP-VAE) - and a second where the encoding distribution is a function of data from both modalities, matching the encoder exactly with that of the MM-GVPAE. In each of these cases, we find that the encoding representation has little effect on model performance. We visualize both the average mean-squared error on held-out data (c) as well as a per-trial mean-squared error scatter-plot (b) to provide a better sense of performance across all models and encoders. We note though that the extent to which MM-GPVAE more accurately identifies latent structure than the unimodal variants is sensitive to the nature of the simulated data and can vary across model fits. For (b) and (c) specifically, we average model performance across 4 random seeds to report overall reconstruction accuracy. We also found through model exploration that GPFA specifically often does comparatively well in identifying latent structure when neural rates are high or have large variance. This is likely true because identifying latents within a linear map is trivial under the standard GPFA model, and so latent accuracy can be very good here if the neural data is sufficiently robust. This means that the metrics reported for the identification of latent structure in Figures 10 and 3 are somewhat dependent on the statistics of the neural rates used to generate and fit the model.
Assessing the importance of pruning the Fourier frequencies
Figure 11 demonstrates the necessity of pruning the Fourier frequencies when evaluating the MM-GPVAE on our simulated data. Here, we present standard and pruned Fourier MM-GPVAE model fits (Fig 11(a)) and show their ability in reconstructing the true generative smooth latent across each latent subspace (Fig 11(b)). Without constraining the number of Fourier features for these data, the latent identification fails, akin to what is seen when we ignore the Fourier representation entirely. We note here that this feature is a function of the length of the trial, and the failure is not as stark for shorter-length trials. For these trials, our pruned condition the length scale of the GP to a value of 10 (i.e. the hyperparameter can go no smaller than 10).
Varying the number of neurons
To demonstrate how the recovered latent of the shared and independent subspaces might trade-off with the amount of data a practitioner may have for a given modality, we consider fits of our MM-GPVAE in our simulated setting with a varying number of neurons. As expected, increasing the number of neurons increases the ability of the MM-GPVAE to identify the neural latent, as well as its ability to accurately reconstruct neural data. Importantly, because structure is shared across modalities, increasing the number of neurons also increases the ability of the MM-GPVAE to identify shared latent structure, as well as shows some improvement in the MM-GPVAE to generate image data. Increasing the number of neurons has no effect on identifying the image latent structure. (See Fig 12)
Robustness to measurement error
To assess how the MM-GPVAE might perform in the presence of noisy or imperfect behavioral measurements, we consider a simulated experiment where a randomly selected of the images are greyed on either the top or bottom half. We fit the MM-GPVAE jointly to the image data of this form as well as uncorrupted neural rates of the same form as in the main manuscript. We find that the MM-GPVAE is able to exploit smoothness, rendering well-generated images even during these occluded trials (Fig 13(a)). As expected, we find that a standard MM-VAE (without a GP prior over time) is not able to reconstruct the true image as well, but does a better job reconstructing the occluded image, as it treats each frame independently (Fig 13(b)). We also show that this occlusion procedure has no bearing on the ability of the MM-GPVAE or MM-VAE to reconstruct the neural rates (Fig 13(c)).
@@
For the MM-GPVAE, both the simulated and real-world multi-modal evaluations used similar neural net architectures. However, there were some modifications of the nodes/layers that were unique to each evaluation. This was necessary as the latent dimensionality was different across our different evaluations, and certain behavioral reconstructions, especially those that were higher-dimensional, required richer neural network parameterizations. The schematic for the MM-GPVAE neural network architecture in our simulated example can be found in Figure 14, and the schematics for the neural network architectures for the real-world multi-modal datasets can be found in Figure 15 (fly) and Figure 16 (hawkmoth). For our evaluation on the dataset used in (Hurwitz et al., 2021), whose results are above, the architecture can be found in Figure 17. Note that for all experiments, each modality is encoded to its own set of variational means and variances (transformed into the Fourier domain). The encoded means and variances representing the shared latents are then summed to give the encoded shared latents means and variances. Across all evaluations, we parameterized our latents in the Fourier domain and converted back to the time domain before decoding.
@@
Pre-processing: We isolated 1000 raw calcium traces from (Schaffer et al., 2021) for use with our model for the evaluations on the Drosophila dataset. For this set-up, we consider a variant of the MM-GPVAE where we have removed the non-linearity for the neural data modality, and instead consider Gaussian observations with an additional parameter controlling the observation variance for the fluorescence traces (akin to the original formulation of GPFA (Yu et al., 2009)). This dataset also contained x,y positions from 8 tracked limbs positions (Mathis et al., 2018). The data came as one continuous recording, and at every time point there was a behavioral categorization probability for one of 6 distinct behaviors (undefined, still, running, front grooming, back grooming, abdomen bending) determined via the algorithm outlined in (Whiteway et al., 2021). To split this continuous recording into trials, we found segments of the recording where, for 35 continuous samples, a single behavior was estimated at a probability. This generated 318 total trials where each trial was of one of 5 possible behaviors. There was no section of the recording where there was an ‘undefined’ behavior for 35 continuous samples, so there are no undefined behavioral trials in our analysis.
Additional evaluation: Though we show the time-course of the behavioral reconstruction in the main manuscript, we show a 2-d reconstruction on 2 example trials in Figure 18. Here we plot true x,y positions for 2 trails for each of the 8 limb positions alongside our model’s reconstruction. We can see here that our model also captures the spatial information in the limb-position data. We also show for comparison the 2-d depiction of the shared and independent latent representation of all trials in the dataset (Figure 19), with all five of the behaviors labelled. You can see here that the "still" behavior separates well from all other behaviors in the shared and neural subspaces.